What property of equality for:
1. If 3a=15, then 15=3a (my answer is symmetric)
2. If b-2=3+4, then b-2=7 (my answer is substitution)

Answer by srinivas.g(540)

If 3a=15, then 15=3a, transitive property

2)If b-2=3+4, then b-2=7 ,substitution property